OpenStudy (anonymous):

How do I evaluate this? 5/2log4 (4^(1/3)

OpenStudy (anonymous):

Do you mean,\[\frac{5}{2}\log_4 4^{1/3}\]?

OpenStudy (anonymous):

If so, then you need to use your log laws:\[\log_4 4^{1/3}=\frac{1}{3}\log _ 4 4=\frac{1}{3}\]

OpenStudy (anonymous):

So your expression is\[\frac{5}{2}\log_4 4^{1/3}=\frac{5}{2}\times \frac{1}{3}=\frac{5}{6}\]